40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 13 03:18:14 2011, kiwamu, Update, LSC, Y arm locking prep image.png
    Reply  Tue Sep 13 19:31:09 2011, kiwamu, Update, LSC, Y arm locking prep Yarmlocked.png
       Reply  Wed Sep 14 01:17:51 2011, kiwamu, Update, LSC, measurement of spot position on Y arm misposition.png
          Reply  Wed Sep 14 15:00:00 2011, kiwamu, Update, LSC, measurement of spot position on Y arm DSC_3408_small.jpg
    Reply  Wed Sep 14 21:18:45 2011, kiwamu, Update, LSC, Y arm locking prep 
Message ID: 5400     Entry time: Wed Sep 14 01:17:51 2011     In reply to: 5398     Reply to this: 5406
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: measurement of spot position on Y arm 

The spot positions on ITMY and ETMY were measured using the LOCKIN modules in C1ASS when the Y arm stayed locked.

The beam was successfully aligned such that it hits the center of the ETMY mirror.

However on the other hand the angle of the beam is pitching and it's going upward as the beam propagates to ETMY.

 

/***** RESULTS ******/

Here is a summary of the measurement :

  amount of off-centering [mm]
ETMY_PIT  0.0032
ETMY_YAW  -0.13
ITMY_PIT  -12
ETMY_YAW  0.91

  

Also a cartoon is shown below.

The scale is not quite true, but at least it gives you a 3D information of how the beam is pointing down to the Y arm.

 misposition.png


  /***** MEASUREMENT *****/

 In order to measure the spot positions the standard technique, namely A2L, was used.

Since the C1ASS model was made for doing the A2L measurements on each arm cavity, the LOCKIN modules in C1ASS were used.

First the Y arm was locked with AS55 (#5398), and then the C1ASS was activated by calling some scripts from C1ASS_QPDs.adl.

In order to calibrate the signals from LOCKINs, an intentional coil imbalance was introduce.

This is the same calibration technique as Valera explained before (#4355) for measurement of the MC spot positions.

Quote from #5398

The Y arm has been locked with AS55.

A next thing is to check the spot positions on the ETMY and ITMY mirrors so that we can evaluate the recent beam pointing.

 

ELOG V3.1.3-